OpenStudy (anonymous):

You wonder how many cars there are in your area. You call up the Ford dealership and ask how many Fords they’ve sold. They report that they’ve sold a total of 15,000 cars. You then take a lawn chair and camp out on a busy road and count cars. You count 1200 cars and note that 700 of them are Fords. Estimate the number of cars in your area

jimthompson5910 (jim_thompson5910):

x = estimated total number of cars \[\Large \frac{\text{number of fords}}{\text{number of cars}}=\frac{\text{number of fords}}{\text{number of cars}}\] \[\Large \frac{15000}{x}=\frac{700}{1200}\] solve for x